OpenStudy (anonymous):

Check whether (– 1, 5) is a solution of each system of equations. x+y=4 x-y=6

OpenStudy (praxer):

I think x has to be equal to 5 and y be -1

OpenStudy (praxer):

as adding both the equations we have 2x=10 =>x=5 and so, y = 4-x=4-5=-1

OpenStudy (anonymous):

b is not statisfing by values if x = -1 and y = 5 but it statisfing by x=5 and y = -1

OpenStudy (praxer):

Which means (-1,5) is not a solution of the system of equations..... :)

OpenStudy (anonymous):

so my answer is not a solution

OpenStudy (praxer):

Yes......... !!!! :)

OpenStudy (anonymous):

thanks